if a cube has a surface area of 294cm2, what are its dimensions and what is its volume?
Found 2 solutions by drj, Alan3354:
Answer by drj(1380)

If a cube has a surface area of 294cm2, what are its dimensions and what is its volume?

Step 1. Surface Area = 6 * Area of Square Face = 6s^2 where s is the side or edge of the cube.

Step 2. Then, 294=6s^2 or 49=s^2 after dividing by 6 to both sides of the equation. Taking the square root to both sides yields s=7.

Step 3. The volume = V = s^3 = 7^3 = 343 cubic cm.

Step 4. ANSWER. The dimensions are 7 cm by 7 cm by 7 cm and the volume is 343 cubic cm.

I hope the above steps were helpful.

For free Step-By-Step Videos on Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra or for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

And good luck in your studies!

Respectfully,
Dr J

Answer by Alan3354(21581)

if a cube has a surface area of 294cm2, what are its dimensions and what is its volume?
---------
A cube has 6 faces, all their areas are equal
294 cm^2/6 = 49 sq cm
----------
Each face is a square, so its area is s^2 (s = length of sides)
s^2 = 49
s = 7 cm
That's its dimensions, all sides are 7 cm
-----------
Vol = s^3 = 7^3
vol = 343 cc